OpenStudy (anonymous):

Find the indicated measure: Diameter, A=48in^2

OpenStudy (anonymous):

Are you sure there isn't a "pi" in the answer?

OpenStudy (anonymous):

since :\[A = \pi r^2 \] then : \[r=\sqrt{A/\pi}\] we have A = 48in^2 , substitute and you will get r = 3.91in r = Diameter/2 which means it's half of the diameter, so D = 2r = 2(3.91) = 7.82 in